# Mechanistic PDE Networks for Discovery of Governing Equations

### v1

@@ Abstract
We present Mechanistic PDE Networks – a model for discovery of governing partial differential equations from data. Mechanistic PDE Networks represent spatiotemporal data as space-time dependent linear partial differential equations in neural network hidden representations. The represented PDEs are then solved and decoded for specific tasks. The learned PDE representations naturally express the spatiotemporal dynamics in data in neural network hidden space, enabling increased power for dynamical modeling. Solving the PDE representations in a compute and memory-efficient way, however, is a significant challenge. We develop a native, GPU-capable, parallel, sparse, and differentiable multigrid solver specialized for linear partial differential equations that acts as a module in Mechanistic PDE Networks. Leveraging the PDE solver, we propose a discovery architecture that can discover nonlinear PDEs in complex settings while also being robust to noise. We validate PDE discovery on a number of PDEs, including reaction-diffusion and Navier-Stokes equations.
@@ 1 Introduction
Partial differential equations (PDEs) enjoy widespread use as interpretable, analytical models of spatio-temporal dynamics in all areas of science Temam (2024); Turing (1952). The development of PDE models still, however, is largely a manual task that is performed by domain experts examining experimental data. The modern explosion of experimental and simulation data has underscored the need for data-driven machine learning models for discovery of governing equations Brunton et al. (2020). In spite of progress Brunton et al. (2016); Raissi et al. (2019); d’Ascoli et al. (2023), the development of machine learning discovery models that are flexible, handle complex nonlinear data and account for missing physics, while being robust to noise and missing data, is very much an open problem.
A recent discovery architecture for ordinary differential equations (ODEs) was demonstrated by Mechanistic Neural Networks (MechNN) Pervez et al. (2024), enabling a flexible and expressive network design for modeling explicit, interpretable nonlinear ODEs. MechNNs embed a specialized, parallel ODE solver in the network, allowing explicit differential equation representations in neural networks, which can be employed for ODE discovery and prediction, both relevant in applications including general time series and numerical simulations Pervez et al. (2024) and in climate science Chen et al. (2025); Yao et al. (2024).
In this paper we focus on governing equation discovery and demonstrate Mechanistic Neural Networks for discovery of partial differential equations from data. We develop a PDE solver, NeuRLP-PDE, for parallel native differentiable solution of linear partial differential equations over multiple spatial dimensions (in addition to the time dimension). In contrast to ODEs, scaling is significantly more difficult in PDE solving, with memory requirements increasing exponentially with dimension and quickly becoming excessive even for mid-sized grids. With NeuRLP-PDE, sparse computation in all phases of the PDE solver ensures that the memory requirement of the solver over batches of PDEs remains limited for sizeable grids. We build the solver using constrained optimization and a sparse multigrid preconditioned iterative linear solver Saad (2003) for solving the large and sparse linear systems that appear in the constrained optimization. The NeuRLP-PDE solver is then incorporated into the PDE discovery architecture enabling the learning of governing PDEs directly from data.
@@ 2 Mechanistic PDE Networks
All functions $u_{i}$ are defined over a domain $\mathcal{D}$ which we assume to be a Cartesian grid, i.e., $\mathcal{D}=\mathcal{T}\times\mathcal{X}_{1}\times\mathcal{X}_{2}\ldots\mathcal{X}_{n_{d}}$, where $\mathcal{T}$, $\mathcal{X}_{i}$ are intervals. For conciseness, we collect the partial derivative indices appearing in PDE $\mathcal{P}_{u_{i}}$ into a multi-index set denoted $\mathcal{M}$. As an example =$\mathcal{M}=\{t,x_{1},x_{2},x_{1}x_{1}\}$, $\{u_{m}\}_{m\in\mathcal{M}}$ denotes the set of partial derivatives $\{u_{t},u_{x_{1}},u_{x_{2}},u_{x_{1}x_{2}}\}$. We also use the empty index $\phi$ for $u$, i.e., $u_{\phi}:=u$. For simplicity of exposition we assume a set with a single PDE denoted $\mathcal{P}_{u}$ for a function $u(t,x)$ over a spatial dimensions $x:=(x_{1},\ldots,x_{n_{d}})$ with $t\in[0,t_{f}]$ and $x_{i}\in[0,x_{i,f}]$.
@@ 3 Linear PDE Solving With Differentiable Optimization
The workhorse of mechanistic PDE networks is NeuRLP-PDE – a specialized, parallel and differentiable solver for linear PDEs. NeuRLP-PDE solves linear PDEs, with partial derivatives of arbitrary order, over a discretized spatio-temporal grid by reducing PDE solution to a relaxed differentiable constrained optimization Young (1961).
@@
Solving these equations gives a solution $z$ of the optimization problem and thus of the PDE. However, we also require the dual variables for the differentiating through the least squares optimization. For this we consider the following saddle-point formulation of least squares Saad (2003).
@@
where $\lambda$ are the dual variables. This formulation can also be seen as the KKT constraints of the convex quadratic programming formulation of least-squares. Given the saddle-point formulation we can apply techniques from differentiable optimization for differentiating through the solution of constrained optimization problems Amos and Kolter (2017).
@@
Backward pass. Give scalar loss $l$ we consider the gradient, $g_{z}:=\partial_{z^{*}}l$, relative to $z^{*}$. We require the gradient of $l$ relative to the matrix $A$ and vector $d$ which contain the optimization data. Following Amos and Kolter (2017), we can achieve this by solving the following saddle-point system:
@@
The gradient of the loss $l$ relative to $A$ and $d$ is $\partial_{A}l=d_{\lambda}{z^{*}}^{\intercal}+\lambda^{*}d_{z}^{\intercal}$ and $\partial_{d}l=-d_{\lambda}$ Amos and Kolter (2017).
Scaling PDE Solving. The linear systems in equations (10), (11), (9) can become very large with increasing dimension of the PDEs which results in very large grids over which the PDE is defined. As an example for a PDE defined over 32x32 grids we have 5120 variables and 9212 constraints requiring 377 MB of double precision dense storage. However for 32x32x32 (220k variables and 420k constraints) we would need 780GB and for 64x64x64 we would need 50k GB. For such large grids the linear systems cannot be solved by dense methods. For the experiments we solve 1D PDEs with a dense solver, since for these for which the memory requirement is reasonably bounded for small grids and batch sizes. However, for 2D or higher PDEs with midsize grids (say, 32 to 64 per dimension), the memory required for dense solvers is too great for GPU storage. The matrices, however, are highly sparse with 99.9% zero values for a 32x32 matrix and 99.999% zeros for a 64x64x64 matrix. For such problems we develop a multigrid preconditioner for use with a sparse iterative solver.
@@ 3.1 An Iterative Sparse Multigrid Solver
To make PDE solving feasible for more than two dimensions (including time) and for large grids, we build an iterative sparse solver for solving the normal equations. We use the FGMRES linear solver, the flexible variant of the Krylov-subspace linear solver GMRES Saad and Schultz (1986), as the base linear solver. As is well-known from numerical linear algebra, standalone iterative solvers can lead to poor solutions for complex problems and require preconditioning for accurate solution. General purpose preconditioning methods such as incomplete LU or Cholesky factorization are not feasible for the large systems we encounter due to their high memory use and lack of GPU implementations. Instead, we build a multigrid V-cycle preconditioner for FGMRES. We briefly describe the multigrid V-cycle here and in Algorithm 1 in the appendix. More details on the multigrid methods and sparse iterative solvers can be found in Saad (2003) and Briggs et al. (2000).
@@ 4 Mechanistic Networks for PDE Discovery
The goal of discovery is to find the equations governing spatio-temporal data representing some physical phenomenon. The result of governing equation discovery is a set of partial differential equations that, ideally, completely describe the given data. Further restrictions may be added to the form of the discovered equations so that the equations may be interpreted by human scientists. One such restriction is that the resulting equations should be composed only of simple functions. Another natural restriction is that the equations should concise, or sparse, with as few terms as possible Brunton et al. (2016).
@@ 5 Related Work
The synthesis of machine learning and differential equations has been treated in a few different ways in the field. A major line of work approaches the problem from a machine learning perspective with supervised or unsupervised learning for partial differential equations Li et al. (2020); Brandstetter et al. (2022). Another line of work in machine learning for scientific application aims at a synthesis of differential equation models with a network-in-solver approach by embedding neural networks in classical solvers Chen et al. (2018); Rackauckas et al. (2020) enabling access to high quality numerical solvers in the context of machine learning. Physics-informed networks use a physics-aware loss to supervise network training Raissi (2018); Raissi et al. (2019). Mechanistic neural networks conversely are a solver-in-network approach that attempts to balance classical solvers and neural networks by combining restricted ODEs Pervez et al. (2024); Chen et al. (2025) (and now PDEs), that allowing faster solving, with neural network learning together with techniques for handling nonlinear equations.
As an application of the synthesis of ML and differential equations various data driven techniques for discovering governing equations have been explored for ODEs such as SINDy Brunton et al. (2016), SINDy-PI Kaheman et al. (2020) and PDEs including PDEFIND Rudy et al. (2017) and weak forms Reinbold et al. (2020); Messenger and Bortz (2021) extending the SINDy framework. Physics informed networks Raissi et al. (2019); Raissi (2018) have also been used for inverse problems but unlike UDEs Rackauckas et al. (2020) and SINDy do not handle unknown physics. PDE-LEARN Stephany and Earls (2024) combines features of physics informed networks and basis libraries for PDE discovery. MechNNs Pervez et al. (2024) build a discovery model for ODEs which we enhance and develop PDE models that handle nonlinear equations and unknown dynamics.
@@ 6 Experiments
Evaluation Metrics. We use two metrics for evaluating the quality of the discovered equation. The true positivity ratio is defined as $\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}+\text{FP}}$ Lagergren et al. (2020) and measures how many terms have been correctly identified. We measure the maximum relative error of the true coefficients as $E_{\infty}(\xi)=\max_{j:\hat{\xi}_{j}^{*}!=0}\frac{|\xi_{j}-\hat{\xi}^{*}_{j}|}{|\xi_{j}|}$, with $\xi,\hat{\xi}$ being the discovered and true coefficients Messenger and Bortz (2021). $E_{\infty}(\xi)$ serves as the measure of accuracy of discovered coefficients.
@@
For 2D PDEs, we work with reaction-diffusion and incompressible Navier-Stokes equations. Ginzburg-Landau reaction-diffusion equations are frequently employed for modeling biological pattern development. We demonstrate discovery in two settings: 1) a relatively easy spiral dataset with a single spiral and 2) a harder setting with significantly more complex patterns where the baseline methods fail to recover the equations even in the absence of noise. We also demonstrate discovery for an incompressible Navier-Stokes example with a small viscosity term Harpole et al. (2019). We set the viscosity term $\nu$ to 0.001 and show that our method is able to recover the equation while the baselines cannot. For baselines, we compare with PDEFIND Rudy et al. (2017) and WeakSINDy Reinbold et al. (2020), with implementations by de Silva et al. (2020), which are known to be robust to noise in many settings. We especially focus on comparing the robustness under noise of our method with the baselines.
@@ 6.1 Solver Unit Test
First, we briefly verify that our solver is capable of solving some given linear PDEs. We test on the 2D Laplace PDE, $\nabla^{2}u=0$ on 32x32, 64x64, and 128x128 grids using the multigrid solver. We compare against a reference finite-difference implementation from the py-pde PyPde package with sinusoidal boundary conditions obtaining $L_{2}$ errors of orders $10^{-1},5\times 10^{-2}$ and $10^{-2}$, respectively, showing the errors decreasing with increasing grid size. This simple evaluation shows that our solver is indeed capable of solving these equations to an accuracy comparable with traditional libraries while also being able to parallelize solving multiple equations on GPUs.
@@ 6.2 PDE Discovery
Easy Case. For the reaction-diffusion equations, we first test on an easy dataset with a single spiral pattern for which the baselines also perform well Reinbold et al. (2020). For this example, the spatial domain has a length of 20 with equal spatial step sizes and a spatial resolution of 64x64 in the $x$ and $y$ dimensions. For time we use 128 time steps with a step size of 0.05. An example of the data at a single time step is shown in Figure 3 (left). The data was generated by a spectral method following the implementation in de Silva et al. (2020), setting $D1=D2=0.1$ and $\beta=1$.
@@
A Harder Case. For reaction diffusion equations we also test on a more difficult dataset with more complex patterns, for which the baselines are unable to discover the correct equations even on noise-free data. The dataset has a spatial resolution of 256x256 over a Cartesian domain of size 10 with equal step size. The data was generated by using the Basilisk solver Kenneally et al. (2020) starting from a random initial condition and running with a maximum time step size of 0.05. We set $D1=D2=1$ and $\beta=1.5$. We take only 128 time steps from the data for training. An example of the data at a single time step is shown in Figure 3 (right).
@@
1D PDEs. We demonstrate discovery on the 1D diffusion equation and inviscid and viscous Burger’s equations. We use degree 4 polynomials for the coefficients of the $u_{x}$ and $u_{xx}$ terms. For all cases in the noise-free case the method discovers the true equation. For Burger’s equation we experiment with noisy data and compare with WeakSINDy and find both methods to be on-par with each other. The results are shown in Appendix A.

### v2

@@ Abstract
We present Mechanistic PDE Networks – a model for discovery of governing partial differential equations from data. Mechanistic PDE Networks represent spatiotemporal data as space-time dependent linear partial differential equations in neural network hidden representations. The represented PDEs are then solved and decoded for specific tasks. The learned PDE representations naturally express the spatiotemporal dynamics in data in neural network hidden space, enabling increased power for dynamical modeling. Solving the PDE representations in a compute and memory-efficient way, however, is a significant challenge. We develop a native, GPU-capable, parallel, sparse, and differentiable multigrid solver specialized for linear partial differential equations that acts as a module in Mechanistic PDE Networks. Leveraging the PDE solver, we propose a discovery architecture that can discover nonlinear PDEs in complex settings while also being robust to noise. We validate PDE discovery on a number of PDEs, including reaction-diffusion and Navier-Stokes equations. Source code will be made available at: https://github.com/alpz/mech-nn-discovery-pde
Machine Learning, ICML
@@ 1 Introduction
Partial differential equations (PDEs) enjoy widespread use as interpretable, analytical models of spatio-temporal dynamics in all areas of science (Temam, 2024; Turing, 1952). The development of PDE models still, however, is largely a manual task that is performed by domain experts examining experimental data. The modern explosion of experimental and simulation data has underscored the need for data-driven machine learning models for discovery of governing equations (Brunton et al., 2020). In spite of progress (Brunton et al., 2016; Raissi et al., 2019; d’Ascoli et al., 2023), the development of machine learning discovery models that are flexible, handle complex nonlinear data and account for missing physics, while being robust to noise and missing data, is very much an open problem.
A recent discovery architecture for ordinary differential equations (ODEs) was demonstrated by Mechanistic Neural Networks (MechNN) (Pervez et al., 2024), enabling a flexible and expressive network design for modeling explicit, interpretable nonlinear ODEs. MechNNs embed a specialized, parallel ODE solver in the network, allowing explicit differential equation representations in neural networks, which can be employed for ODE discovery and prediction, both relevant in applications including general time series and numerical simulations (Pervez et al., 2024) and in climate science (Chen et al., 2025; Yao et al., 2024).
In this paper we focus on governing equation discovery and demonstrate Mechanistic Neural Networks for discovery of partial differential equations from data. We develop a PDE solver, NeuRLP-PDE, for parallel native differentiable solution of linear partial differential equations over multiple spatial dimensions (in addition to the time dimension). In contrast to ODEs, scaling is significantly more difficult in PDE solving, with memory requirements increasing exponentially with dimension and quickly becoming excessive even for mid-sized grids. With NeuRLP-PDE, sparse computation in all phases of the PDE solver ensures that the memory requirement of the solver over batches of PDEs remains limited for sizeable grids. We build the solver using constrained optimization and a sparse multigrid preconditioned iterative linear solver (Saad, 2003) for solving the large and sparse linear systems that appear in the constrained optimization. The NeuRLP-PDE solver is then incorporated into the PDE discovery architecture enabling the learning of governing PDEs directly from data.
@@ 2 Mechanistic PDE Networks
All functions $u_{i}$ are defined over a domain $\mathcal{D}$ which we assume to be a Cartesian grid, i.e., $\mathcal{D}=\mathcal{T}\times\mathcal{X}_{1}\times\mathcal{X}_{2}\ldots\mathcal{X}_{n_{d}}$, where $\mathcal{T}$, $\mathcal{X}_{i}$ are intervals. For conciseness we collect the partial derivative indices appearing in PDE $\mathcal{P}_{u_{i}}$ into a multi-index set denoted $\mathcal{M}$. As an example =$\mathcal{M}=\{t,x_{1},x_{2},x_{1}x_{2}\}$, $\{u_{m}\}_{m\in\mathcal{M}}$ denotes the set of partial derivatives $\{u_{t},u_{x_{1}},u_{x_{2}},u_{x_{1}x_{2}}\}$. We also use the empty index $\phi$ for $u$, i.e., $u_{\phi}:=u$. For simplicity of exposition we assume a set with a single PDE denoted $\mathcal{P}_{u}$ for a function $u(t,x)$ over a spatial dimensions $x:=(x_{1},\ldots,x_{n_{d}})$ with $t\in[0,t_{f}]$ and $x_{i}\in[0,x_{i,f}]$.
@@ 3 Linear PDE Solving With Differentiable Optimization
The workhorse of mechanistic PDE networks is NeuRLP-PDE – a specialized, parallel and differentiable solver for linear PDEs. NeuRLP-PDE solves linear PDEs, with partial derivatives of arbitrary order, over a discretized spatio-temporal grid by reducing PDE solution to a relaxed differentiable constrained optimization (Young, 1961).
@@
Solving these equations gives a solution $z$ of the optimization problem and thus of the PDE. However, we also require the dual variables for the differentiating through the least squares optimization. For this we consider the following saddle-point formulation of least squares (Saad, 2003).
@@
where $\lambda$ are the dual variables. This formulation can also be seen as the KKT constraints of the convex quadratic programming formulation of least-squares. Given the saddle-point formulation we can apply techniques from differentiable optimization for differentiating through the solution of constrained optimization problems (Amos & Kolter, 2017).
@@
Backward pass. Give scalar loss $l$ we consider the gradient, $g_{z}:=\partial_{z^{*}}l$, relative to $z^{*}$. We require the gradient of $l$ relative to the matrix $A$ and vector $d$ which contain the optimization data. Following Amos & Kolter (2017), we can achieve this by solving the following saddle-point system:
@@
The gradient of the loss $l$ relative to $A$ and $d$ is $\partial_{A}l=d_{\lambda}{z^{*}}^{\intercal}+\lambda^{*}d_{z}^{\intercal}$ and $\partial_{d}l=-d_{\lambda}$ (Amos & Kolter, 2017).
Scaling PDE Solving. The linear systems in equations (10), (11), (9) can become very large with increasing dimension of the PDEs resulting in very large grids over which the PDE is defined. As an example, for a PDE defined over 32x32 grids we have 5120 variables and 9212 constraints, requiring 377 MB of double precision dense storage. However for 32x32x32 (220k variables and 420k constraints) we would need 780GB and for 64x64x64 we would need 50k GB. For such large grids the linear systems cannot be solved by dense methods. For the experiments we solve 1D PDEs with a dense solver since the memory requirement is reasonably bounded for small grids and batch sizes. However, for 2D or higher PDEs with midsize grids (say, 32 to 64 per dimension), the memory required for dense solvers is too great for GPU storage. The matrices, however, are highly sparse with 99.9% zero values for a 32x32 matrix and 99.999% zeros for a 64x64x64 matrix. For such problems we develop a multigrid preconditioner for use with a sparse iterative solver.
@@ 3.1 An Iterative Sparse Multigrid Solver
To make PDE solving feasible for more than two dimensions (including time) and for large grids, we build an iterative sparse solver for solving the normal equations. We use the FGMRES linear solver, the flexible variant of the Krylov-subspace linear solver GMRES (Saad & Schultz, 1986), as the base linear solver. As is well-known from numerical linear algebra, standalone iterative solvers can lead to poor solutions for complex problems and require preconditioning for accurate solution. General purpose preconditioning methods such as incomplete LU or Cholesky factorization are not feasible for the large systems we encounter due to their high memory use and lack of GPU implementations. Instead, we build a multigrid V-cycle preconditioner for FGMRES. We briefly describe the multigrid V-cycle here and in Algorithm 1 in the appendix. More details on the multigrid methods and sparse iterative solvers can be found in Saad (2003) and Briggs et al. (2000).
@@ 4 Mechanistic Networks for PDE Discovery
The goal of discovery is to find the equations governing spatio-temporal data representing some physical phenomenon. The result of governing equation discovery is a set of partial differential equations that, ideally, completely describe the given data. Further restrictions may be added to the form of the discovered equations so that the equations may be interpreted by human scientists. One such restriction is that the resulting equations should be composed only of simple functions. Another natural restriction is that the equations should concise, or sparse, with as few terms as possible (Brunton et al., 2016).
@@ 5 Related Work
The synthesis of machine learning and differential equations has been treated in a few different ways in the field. A major line of work approaches the problem from a machine learning perspective with supervised or unsupervised learning for partial differential equations (Li et al., 2020; Brandstetter et al., 2022). Another line of work in machine learning for scientific application aims at a synthesis of differential equation models with a network-in-solver approach by embedding neural networks in classical solvers (Chen et al., 2018; Rackauckas et al., 2020) enabling access to high quality numerical solvers in the context of machine learning. Physics-informed networks use a physics-aware loss to supervise network training (Raissi, 2018; Raissi et al., 2019). Mechanistic neural networks conversely are a solver-in-network approach that attempts to balance classical solvers and neural networks by combining restricted ODEs (Pervez et al., 2024; Chen et al., 2025) (and now PDEs), that allowing faster solving, with neural network learning together with techniques for handling nonlinear equations.
As an application of the synthesis of ML and differential equations various data driven techniques for discovering governing equations have been explored for ODEs such as SINDy (Brunton et al., 2016), SINDy-PI (Kaheman et al., 2020) and PDEs including PDEFIND (Rudy et al., 2017) and weak forms (Reinbold et al., 2020; Messenger & Bortz, 2021) extending the SINDy framework. Physics informed networks (Raissi et al., 2019; Raissi, 2018) have also been used for inverse problems but unlike UDEs (Rackauckas et al., 2020) and SINDy do not handle unknown physics. PDE-LEARN (Stephany & Earls, 2024) combines features of physics informed networks and basis libraries for PDE discovery. MechNNs (Pervez et al., 2024) build a discovery model for ODEs which we enhance and develop PDE models that handle nonlinear equations and unknown dynamics.
@@ 6 Experiments
Evaluation Metrics. We use two metrics for evaluating the quality of the discovered equation. The true positivity ratio is defined as $\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}+\text{FP}}$ (Lagergren et al., 2020) and measures how many terms have been correctly identified. We measure the maximum relative error of the true coefficients as $E_{\infty}(\xi)=\max_{j:\hat{\xi}_{j}^{*}!=0}\frac{|\xi_{j}-\hat{\xi}^{*}_{j}|}{|\xi_{j}|}$, with $\xi,\hat{\xi}$ being the discovered and true coefficients (Messenger & Bortz, 2021). $E_{\infty}(\xi)$ serves as the measure of accuracy of discovered coefficients.
@@
For 2D PDEs, we work with reaction-diffusion and incompressible Navier-Stokes equations. Ginzburg-Landau reaction-diffusion equations are frequently employed for modeling biological pattern development. We demonstrate discovery in two settings: 1) a relatively easy spiral dataset with a single spiral and 2) a harder setting with significantly more complex patterns where the baseline methods fail to recover the equations even in the absence of noise. We also demonstrate discovery for an incompressible Navier-Stokes example with a small viscosity term (Harpole et al., 2019). We set the viscosity term $\nu$ to 0.001 and show that our method is able to recover the equation while the baselines cannot. For baselines, we compare with PDEFIND (Rudy et al., 2017) and WeakSINDy (Reinbold et al., 2020), with implementations by de Silva et al. (2020), which are known to be robust to noise in many settings. We especially focus on comparing the robustness under noise of our method with the baselines.
@@ 6.1 Solver Unit Test
First, we verify that our solver is capable of solving some given linear PDEs. We test on the 2D Laplace PDE, $\nabla^{2}u=0$ on 32x32, 64x64, and 128x128 grids using the multigrid solver. We compare against a reference finite-difference implementation from the py-pde (PyPde, ) package with sinusoidal boundary conditions obtaining $L_{2}$ errors of orders $10^{-1},5\times 10^{-2}$ and $10^{-2}$, respectively, showing the errors decreasing with increasing resolution. This simple evaluation shows that our solver is indeed capable of solving these equations to an accuracy comparable with traditional libraries while also being able to parallelize solving multiple equations on GPUs.
@@ 6.2 PDE Discovery
Easy Case. For the reaction-diffusion equations, we first test on an easy dataset with a single spiral pattern for which the baselines also perform well (Reinbold et al., 2020). For this example, the spatial domain has a length of 20 with equal spatial step sizes and a spatial resolution of 64x64 in the $x$ and $y$ dimensions. For time we use 128 time steps with a step size of 0.05. An example of the data at a single time step is shown in Figure 3 (left). The data was generated by a spectral method following the implementation in de Silva et al. (2020), setting $D1=D2=0.1$ and $\beta=1$.
@@
A Harder Case. For reaction diffusion equations we also test on a more difficult dataset with more complex patterns, for which the baselines are unable to discover the correct equations even on noise-free data. The dataset has a spatial resolution of 256x256 over a Cartesian domain of size 10 with equal step size. The data was generated by using the Basilisk solver (Kenneally et al., 2020) starting from a random initial condition and running with a maximum time step size of 0.05. We set $D1=D2=1$ and $\beta=1.5$. We take only 128 time steps from the data for training. An example of the data at a single time step is shown in Figure 3 (right).
@@
1D PDEs. We demonstrate discovery on the 1D diffusion equation and inviscid and viscous Burger’s equations. We use degree 4 polynomials for the coefficients of the $u_{x}$ and $u_{xx}$ terms. For all cases in the noise-free case the method discovers the true equation. For Burger’s equation we experiment with noisy data and compare with WeakSINDy and find both methods to be on-par with each other. The results for the are shown in Figure 7 and Appendix A.